OpenStudy (anonymous):

Can someone please help me I have to prove this equation is true using algebraic proof, identifying each step, and demonstrate that it works on numerical relationships. The equation is (x-y)(x-y)=x^2-xy-yx+y. All I have so far is (2-3)(2-3)=2^2-2x3-3x2+3. Help would be most appreciated.

OpenStudy (anonymous):

(x-y)(x-y)=x^2-2(xy)+y^2 would be the correct way. you could show that using the distributive property and it would explain it (the distributive property is when you multiply all variables or terms within a set of parentheses by a constant, variable, or all terms in another set of parentheses).

OpenStudy (anonymous):

multiply (second bracket ) by each TERM in the( first bracket) x*x +x*(-y) then -y *x +(-y)*(-y) x^2 -xy -xy +y^2 because miinus times minus gives positive I notice you did not have y^2 in your question a typo?? so not 3 but 3^2 =9 so you have 4-6-6+9=1 (2-3)(2-3) (-1)(-1) =+1 as your proof